plm for each time step. Use the approximation that aye A paddle ball toy consists of a flat wooden paddle and a small rubberPart 2 (b) Starting with the same initial position (-0.2, -0.61, 0> m) and momentum (-0.02, -o.01, -0.02 > kg m/s) determinePlease help wtih all 3!
Answer #1

mass of the ball m = 0.015 kg

string length l = 0.330 m

elastic const. k = 0.95 N/m

momentum of the ball  p = <-0.02,-0.01,-0.02> kg-m/s

position  r = <-0.2,-0.61,0> m

\Deltat =0.1 s

average velocities of the ball

vx = -0.02 m/s ; vy = -0.01 m/s ; vz = -0.02 m/s

after 0.1s

x= -0.2 -0.02*0.1 = -0.202 m

y = -0.61 - 0.01*0.1 = -0.611 m

z= 0 -0.02*0.1 = -0.002 m

position at 0.1 s r(0.1) = <-0.202, -0.611, -0.002>

c)

average velocity \Delta r/\Deltat = p/m

We only know the final momentum p , we do not know the change in momentum \Delta p and the \Delta t, time interval. We cannot estimate the force and the acceleration. We can only compute the average velocity from the final momentum. Using average velocity, the motion is uniform and it will not make any difference of the size in interval \Delta t , if it is 0.05 s or 0.1 s

The answer will not be different and any arbitrary time step will do.